OpenStudy (anonymous):

Please help I Will fan and give medals Have multiple Questions!!!!!!! Need a smarty(: Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is zero undefined positive negative

OpenStudy (a1234):

 Can you picture which one it is from the points?

OpenStudy (anonymous):

Not sure @a1234

OpenStudy (a1234):

 Connect the two points. What do you get?

OpenStudy (anonymous):

@a1234 Would it be positive?

OpenStudy (a1234):

Good job! It's positive.

OpenStudy (anonymous):

@a1234 Thank you very much for your help! anyway you could help me with a few more?

OpenStudy (a1234):

Yes, I can.

OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2). y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

@a1234 Having trouble here I Got B as my answer....

OpenStudy (a1234):

Use y - y1 = m(x - x1). You have y = 2x + 4 and the point (3, -2). Plug the values in accordingly. y + 2 = 2(x - 3) y + 2 = 2x - 6 y = 2x - 8

OpenStudy (anonymous):

So having y = 2x - 8 as the answer???? @a1234

OpenStudy (a1234):

Yes.

OpenStudy (anonymous):

@a1234 Thank you so much your awesome i have 2 more if you wouldnt mind...... The equation of line AB is (y−3) = 5 (x − 4). What is the slope of a line perpendicular to line AB? −5 5 negative 1 over 5 The equation of line QR is y = negative 1 over 2x + 1. Write an equation of a line perpendicular to line QR in slope-intercept form that contains point (5, 6). y = 2x + 16 y = negative 1 over 2x + 17 over 2 y = − 1 over 2x + 7 over 2 y = 2x − 4 1 over 5

OpenStudy (a1234):

The slopes of perpendicular lines equal -1. Expand (y−3) = 5 (x − 4). y - 3 = 5x - 20 y = 5x - 17 The slope of this line is 5. So, if you need a product of -1, then the other slope must be -1/5. y - y1 = m(x - x1) y + 3 = -1/5(x - 4) y + 3 = -1/5x + 4/5 y = -1.5x - 2.2 So, do you see what the slope is?

OpenStudy (anonymous):

@a1234 The first question being -1/5?????? And the second one being?

OpenStudy (anonymous):

@a1234 B?

OpenStudy (a1234):

The second one...you give it a try :). Why do you say B?

OpenStudy (a1234):

Wait... what exactly does negative 1 "over" mean?

OpenStudy (anonymous):

@a1234 Im not understanding ...?

OpenStudy (a1234):

What does "-1 over 2x + 1" mean in the second problem? You try solving the problem this time.

OpenStudy (anonymous):

@a1234 i told You what i thought was the answer...? Can you confirm the answer???

OpenStudy (a1234):

I don't even exactly know what the question means. How can I confirm the answer?

OpenStudy (anonymous):

The equation of line QR is y = -1/2x + 1. Write an equation of a line perpendicular to line QR in slope-intercept form that contains point (5, 6). y = 2x + 16 y = negative 1 over 2x + 17 over 2 y = − 1 over 2x + 7 over 2 y = 2x − 4

OpenStudy (anonymous):

@a1234 Better??? sorry about that

OpenStudy (a1234):

y = negative 1 over 2x + 1 What does this mean?

OpenStudy (anonymous):

QR is y = -1/2x + 1 @a1234

OpenStudy (a1234):

Aaah...much better :). -1/2 times what number will equal -1?

OpenStudy (anonymous):

@a1234 Yes hahhaha............ .......... Would it be postive 1/2?

OpenStudy (a1234):

No, -1/2 * 1/2 = -1/4

OpenStudy (anonymous):

@a1234 Oh thetn im not sure

OpenStudy (a1234):

That number is just the opposite reciprocal, which is 2 in this case. You have y = -1/2x + 1 and the point (5, 6). Use y - y1 = m(x - x1) and plug in the points like I did in the previous problem. What do you get?

OpenStudy (a1234):

y - 6 = 2(x - 5) y - 6 = 2x - 10 y = 2x - 4 So, yes, y = 2x - 4 :).

OpenStudy (anonymous):

@a1234 Thank you so much for all your help your amazing!!!!!!!!!!!!!!!!!!!1

OpenStudy (a1234):

You're welcome.